KV-COHOMOLOGY OF CONTACT MANIFOLDS

Michel Nguiffo BOYOM

Abstract

Given a contact manifold (M, a) the group of a-preserving diffeomorphisms is denoted by G(a).We construct a Koszul-Vinberg chain complex C(a) on which G(a) acts by chaincomplex homomorphism. The G(a)-equivariant cohmology spaces of C(a) produce new contact invariants.

NGUIFFO BOYOM, Michel Departement de Mathématiques Université Montpellier2, FRANCE e-mail: boyom@math.univ-montp2.fr